OpenStudy (anonymous):

if f(x) = 1/x^2 find f''(x)

OpenStudy (anonymous):

f'(x) = -2/x^3 f''(x) = 6/x^4

OpenStudy (loser66):

why don't you close it?

OpenStudy (anonymous):

close what?

OpenStudy (loser66):

this post

OpenStudy (loser66):

I will take care of the answers, if you see the helper have medal, that means the answer is correct. then close and post another one, ok?